If a+b=11and a 2 +b 2 =65, then what is the value of a 3 +b 3 ?

If a+b=11and a2+b2=65, then what is the value of a3+b3?

Grade:9

2 Answers

Arun
6 years ago
a+b = 11 and,a²+b²=65 so ,(a+b)² = a²+b²+2ab , Put 11 in a+b and 65 in a²+b² to get ab(11)²= 65 + 2ab121 = 65 + 2ab2ab = 56ab = 28Put the value of ab in :a³+b³ = (a+b)³ - 3ab( a+b) = (11)³ - 3×28(11) = 1331 - 924 = 407
Sidhant
6 years ago
A + b is equal to 11 and a square + b square is equal to 65 so a + b square is equal to a square + b square + 280 foot 11 in a + b and 65 in a square + b square to get a b 11 square is equal to 65 + 28121 is equal to 65 + b square is equal to 56 KB is equal to 28 foot with the value of a b in the ratio a cube plus b cube is equal to a + b cube create a + b is equal to 3 minus 3 = 28 in 11 is equal to 1331 - 924 =407